OpenStudy (anonymous):

Find the sum of the first seven terms of the sequence 4, 12, 36, 108, … .

OpenStudy (anonymous):

any one please help???

OpenStudy (anonymous):

Have you read geometric progressions yet ???

OpenStudy (anonymous):

no

OpenStudy (anonymous):

1,456 1,457 4,372 4,374

OpenStudy (anonymous):

can you list out the first 7 terms of the series ??

OpenStudy (anonymous):

it only gives me 4

OpenStudy (anonymous):

4,12,36,108

OpenStudy (anonymous):

look at the series... 12/4 = 3 36/12 = 3 108/36 = 3 this kind of sequence is called a geometric progression with first term "a"=4 and common ratio "r"=3

OpenStudy (anonymous):

ahh so i would multiply to find the rest?

OpenStudy (anonymous):

yup...

OpenStudy (anonymous):

sweet thanks!!!!

OpenStudy (anonymous):

you are welcome